If A and Bare vector fields, prove the following:



nabla(A* B)=(B* nabla)A+(A* nabla)B+B*( nabla* A)+A*( nabla* B) .

1
Expert's answer
2021-12-07T07:03:34-0500

Consider "\nabla(\vec{A} \cdot \vec{B})=\sum \vec{i} \frac{\partial}{\partial x}(\vec{A} \cdot \vec{B})=\sum \vec{i}\left(\frac{\partial \vec{A}}{\partial x} \cdot \vec{B}+\vec{A} \cdot \frac{\partial \vec{B}}{\partial x}\right)"

"=\sum \vec{i}\left(\frac{\partial \vec{A}}{\partial x} \cdot \vec{B}\right)+\sum_{i}\left(\vec{A}, \frac{\partial \vec{B}}{\partial x}\right) \ldots \ldots(i)"

Now "\vec{B} \times\left(\vec{i} \times \frac{\partial \vec{A}}{\partial x}\right)=\left(\vec{B} \cdot \frac{\partial \vec{A}}{\partial x}\right) \vec{i}-(\vec{B}, \vec{i}) \frac{\partial \vec{A}}{\partial x}"

"\Rightarrow\left(\vec{B} \cdot \frac{\partial \vec{A}}{\partial x}\right) \vec{i}=\vec{B} \times\left(\vec{i} \times \frac{\partial \vec{A}}{\partial x}\right)+(\vec{B} \cdot \vec{i}) \frac{\partial \vec{A}}{\partial x}"

"\therefore \sum\left(\vec{B} \cdot \frac{\partial \vec{A}}{\partial x}\right) \vec{i}=\sum \vec{B} \times\left(\vec{i} \times \frac{\partial \vec{A}}{\partial x}\right)+\sum(\vec{B} \cdot \vec{i}) \frac{\partial \vec{A}}{\partial x}"

"=\vec{B} \times \sum\left(\vec{i} \times \frac{\partial \vec{A}}{\partial x}\right)+\left(\vec{B} \cdot \sum \vec{i} \frac{\partial}{\partial x}\right) \vec{A}"

"=\vec{B} \times(\nabla \times \vec{A})+(\vec{B} \cdot \nabla) \vec{A}"

"\therefore \sum \vec{i}\left(\frac{\partial \vec{A}}{\partial x} \vec{B}\right)=\vec{B} \times(\nabla \times \vec{A})+(\vec{B} \cdot \nabla) \vec{A}"

Similarly, by interchanging the roles of "\vec{A}" and "\vec{B}" , we can prove

 "\sum \vec{i}\left(\vec{A} \cdot \frac{\partial \vec{B}}{\partial x}\right)=\vec{A} \times(\nabla \times \vec{B})+(\vec{A} \cdot \nabla) \vec{B}"  

Substituting (ii) and (iii) in (i), we get

 "\nabla(\vec{A} \cdot \vec{B})=\vec{B} \times(\nabla \times \vec{A})+(\vec{B} \cdot \nabla) \vec{A}+\vec{A} \times(\nabla \times \vec{B})+(\vec{A} \cdot \nabla) \vec{B}"